x^2+y^2-2xy=1
3x-y=5

Answer by mananth(16949)

x^2+y^2-2xy=1
3x-y=5
y=3x-5
substitute y in x^2+y^2-2xy=1
x^2+(3x-5)^2-2*x*(3x-5)=1
x^2+9x^2-30x+25-6x^2+10x=1
4x^2-20x+25=1
4x^2-20x+24=0
/4
x^2-5x+6=0
x^2-3x-2x+6=0
x(x-3)-2(x-3)=0
(x-3)(x-2)=0
x= 2 OR 3
Plug x y=3x-5
When x=2, y = 1 (2,1)
when x=3,y=4. (3,4)
m.ananth@hotmail.ca